OpenStudy (anonymous):

a cellular phone plan had the following schedule:

OpenStudy (anonymous):

Please give more information...

OpenStudy (anonymous):

basic service, including 100 mins of calls= $20 per month 2nd 100 mins of calls= .075 per month additional mins= .10 per month construct a function that relates the monthly charge C for x mins of calls

OpenStudy (anonymous):

It's going to need to be piece wise. Fill in the blanks.\[C(x)=\left\{\begin{array}{ll}\cdots & \text{if $x \le 100$} \\ \cdots & \text{if $100 < x \le 200$} \\ \cdots & \text{if $x > 200$}\end{array}\right.\]

OpenStudy (anonymous):

thank you